Please help solve Thanks
Each question on a five-question multiple choice quiz has answer choices labeled A,B,C,and D. How many different ways can a student answer the five questions?
Is the answer 5*4=20 or 5P4=120 or ????

Answer by sudhanshu_kmr(1152)


A student can answer a question in 4 ways.
Total number of ways to answer 5 questions = 4 * 4 * 4 * 4 * 4 = 4^5 = 1024

If any doubt, feel free to contact by mail...